brad can paint a fence in 8 hours. ashley can do it in 4 hours. how long will it take them to do the job working together

Answer by checkley79(3341)

1/8+1/4=1/X FIND THE LCD. (8)
(1+2*1)/8=1/X
(1+2)/8=1/X
3/8=1/X CROSS MULTIPLY.
3X=8
X=8/3
X=2.67 HOURS. ANS.